a certain type of battery has a mean shelf life of 600 days with a standard deviation of 28 days


1
Expert's answer
2021-02-28T15:03:40-0500

Q:What is the probability that the shelf life of the battery is over 630 days?

Mean("\mu" )= 600

Standard deviation("\sigma" )= 28


Normal distribution formula"\Rightarrow z=\dfrac{\bar x-\mu}{\sigma}"


"\Rightarrow z=\dfrac{630-600}{28}"

"\Rightarrow z=1.071"

"P(z>1.071)= 0.5-0.3577=0.1423"